# Unsupervised Machine Learning for Classifying CHIME Fast Radio Bursts and Investigating Empirical Relations

### v1

@@ Abstract
Fast Radio Bursts (FRBs) are highly energetic millisecond-duration astrophysical phenomena typically categorized as repeaters or non-repeaters. However, observational limitations may lead to misclassifications, suggesting a larger proportion of repeaters than currently identified. In this study, we leverage unsupervised machine learning techniques to classify FRBs using data from the CHIME/FRB catalog, including both the first catalog and a recent repeater catalog. By employing Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction and clustering algorithms (k-means and HDBSCAN), we successfully segregate repeaters and non-repeaters into distinct clusters, identifying over 100 potential repeater candidates. Our analysis reveals several empirical relations within the clusters, including the ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$, ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$, and $r-\gamma$ correlations, which provide new insights into the physical properties and emission mechanisms of FRBs. This study demonstrates the effectiveness of unsupervised learning in classifying FRBs and identifying potential repeaters, paving the way for more precise investigations into their origins and applications in cosmology. Future improvements in observational data and machine learning methodologies are expected to further enhance our understanding of FRBs.
@@ 1 Introduction
To explore their origin, numerous radio telescopes have been constructed, such as, the Deep Synoptic Array (DSA; J. et al., 2019; Hallinan et al., 2019), Arecibo (Spitler et al., 2014), Parkes (e.g., Lorimer et al., 2007; Burke-Spolaor & Bannister, 2014; Petroff et al., 2015; Ravi & Lasky, 2014), the Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME/FRB Collaboration et al., 2018), the Five-hundred-meter Aperture Spherical Radio Telescope (FAST; Li & Pan, 2016), and the Australian Square Kilometer Array Pathfinder (ASKAP; Shannon et al., 2024). So far, nearly a thousand FRBs have been observed(Petroff et al., 2016; CHIME/FRB Collaboration et al., 2021; Jankowski et al., 2023; Xu et al., 2023), with nearly 100 of them have known redshifts (e.g. Law et al., 2024; Sharma et al., 2024; Bhardwaj et al., 2024; Gordon et al., 2023). To date, the largest FRB sample is the first CHIME/FRB catalog (CHIME/FRB Collaboration et al., 2021).
In astronomical observations, the measured quantities are often influenced by the distance between the source and Earth. Therefore, distance-related information is crucial for analyzing the origin of FRBs. A key observational parameter correlated with distance for FRBs is the Dispersion Measure (DM), which quantifies the total column density of free electrons along the line of sight between the source and the observer. For the vast majority of observed FRBs, the DM values significantly exceed the predicted values for their respective directions within the Milky Way, indicating that these FRBs originate from distant extragalactic regions of the universe. Meanwhile, an FRB event, FRB 20200428, was observed to have originated within the Milky Way, providing evidence that at least some FRBs are associated with magnetars (Andersen et al., 2020; Li et al., 2021a; Bochenek et al., 2020; Lin et al., 2020).
@@
Based on the observational characteristics, FRBs are broadly classified into two categories, i.e., repeaters and non-repeaters. As the names suggest, repeaters are sources that have exhibited multiple bursts, whereas non-repeaters have only been observed once. Since the discovery of the first repeater, FRB 20121102 (Spitler et al., 2016), more than 50 FRBs have been identified as repeaters to date (CHIME/FRB Collaboration et al., 2021, 2023; Kumar et al., 2019; Kirsten et al., 2022; Fonseca et al., 2020; Niu et al., 2022; Xu et al., 2022), while most of the remaining ones are classified as non-repeaters. However, Ravi (2019) suggests that the volumetric rate of non-repeating FRBs might exceed that of cataclysmic events and the formation rate of compact objects, implying that the majority of FRBs should be repeaters. Some studies also propose that more than half of the FRBs in the first CHIME/FRB catalog could be repeaters (Yamasaki et al., 2023; McGregor & Lorimer, 2024). Furthermore, several FRBs initially identified as non-repeaters were later observed to show repeating characteristics (CHIME/FRB Collaboration et al., 2023). This suggests that many of the FRBs currently classified as non-repeaters might actually be potential repeaters, with only a single burst detected due to various observational factors. As a result, some studies have attempted to identify potential repeaters among apparent non-repeating FRBs, and one of the methods being employed is machine learning.
@@
Machine learning is an artificial intelligence technique that allows computers to learn from data and make predictions or decisions without being explicitly programmed (Cover & Hart, 1967; Rumelhart et al., 1986). Machine learning is generally categorized into three types, i.e., supervised learning, unsupervised learning, and semi-supervised learning (e.g. Dempster et al., 2018; Breiman, 2001; Chang & Lin, 2011; Vapnik, 1999). Using algorithms to uncover patterns in the data can be applied to tasks such as classification, regression, clustering, and optimization. Until now, machine learning has already been widely used in the detection and analysis of FRBs (e.g. Wagstaff et al., 2016; Zhang et al., 2018; Wu et al., 2019; Yang et al., 2021; Adámek & Armour, 2020; Agarwal et al., 2020; Bhatporia et al., 2023). For instance within the first CHIME/FRB catalog dataset, Chen et al. (2021) and Zhu-Ge et al. (2022) identify 188 and 117 repeater candidates from 474 non-repeating FRBs via unsupervised learning, correspondingly. Yang et al. (2023) instead discovered 145 repeater using FRB morphology as features, while Luo et al. (2022) identified dozens of repeater candidates with varies supervised learning methods. Furthermore, Luo et al. (2022) found that the most prominent factors to distinguish between non-repeating and repeating FRBs are brightness temperature and rest-frame frequency bandwidth, whereas Sun et al. (2024) found spectral running may play a role instead. Furthermore, some studies have used machine learning techniques to classify thousands of bursts from highly active repeaters, such as FRB 20121102 (Raquel et al., 2023) and FRB 20201124A (Chen et al., 2023), in an effort to analyze their potential radiation mechanisms.
@@
In addition to classifying FRBs based on their repeatability, some studies have also investigated the clustering of FRBs using characteristics beyond repeatability, as well as the empirical relationships among these features. For instance, similar to the well-known classification of GRB, where short GRBs are short lived and associated with old population, while long GRB with long duration and associated with young population (Zhang et al., 2007; Kumar & Zhang, 2014), Guo & Wei (2022) proposed classifying FRBs based on their association with either old or young stellar populations. They discovered several tight empirical relations for non-repeaters in the first CHIME/FRB catalog, such as ${\rm log\,}E-{\rm log\,}L_{\nu}$, ${\rm log\,}E-{\rm log\,DM_{E}}$ and ${\rm log\,DM_{E}}-{\rm log\,}L_{\nu}$, where ${\rm log\,}$ means the logarithm to base 10, and $E$, ${\rm DM_{E}}$, $L_{\nu}$ are isotropic energy, spectral luminosity, and extragalactic DM, respectively. Similar empirical relations were found for non-repeaters associated with old populations and all non-repeaters, though with notably different slopes and intercepts, such as ${\rm log\,}F_{\nu}-{\rm log\,}S_{\nu}$, ${\rm log\,DM_{E}}-{\rm log\,}S_{\nu}$, and ${\rm log\,}F_{\nu}-{\rm log\,DM_{E}}$, where $F_{\nu}$ is specific fluence, $S_{\nu}$ is flux. Many empirical relations still hold for localized FRBs (Li et al., 2024). Based on these empirical relations, FRBs could potentially serve as standard candles, allowing cosmological models to be constrained without relying on DM measurements (Guo & Wei, 2024). Li et al. (2021b) classified FRBs into short ($W<100\,$ms) and long ($W>100\,$ms) bursts, where $W$ represents the pulse width. A strong power-law correlation between fluence and peak flux density was identified for these categories. Xiao & Dai (2022) classified repeating FRBs into classical ($T_{B}\geq 10^{33}\,$) and atypical ($T_{B}<10^{33}\,$) bursts, where $T_{B}$ refers to the brightness temperature. A tight power-law correlation between pulse width and fluence was also observed for classical bursts.
@@
CHIME has recently released a new repeater catalog (CHIME/FRB Collaboration et al., 2023), significantly increasing the data available on repeaters. This expanded dataset is expected to improve the accuracy of identifying repeater candidates through machine learning. In this paper, we will apply unsupervised learning methods to classify the FRBs from the combined data of these two catalogs into different clusters, find potential repeaters among the non-repeaters, and analyze possible empirical relations across these clusters. This paper is structured as follows. In Section 2, we present the selected CHIME data and the features used in our analysis. introduces two types of unsupervised machine learning methods, i.e., dimensionality reduction and clustering, as well as evaluation metrics. In Section 4, we present the results of dimensionality reduction and clustering, then we analyze the potential empirical relations. Finally, Section 5 provides our conclusions.
@@ 2.1 Sample Construction
In this paper, we use the largest available FRB data from CHIME to date, including the first CHIME/FRB Catalog (CHIME/FRB Collaboration et al., 2021, here after Cat1) and the CHIME/FRB Collaboration (2023) Catalog (CHIME/FRB Collaboration et al., 2023, here after Cat2023). The Cat1 including 536 events (474 non-repeaters and 62 repeat bursts from 18 repeaters), all events have 600 sub-bursts because of multiple peaks appearing in the light curve of FRB. The Cat2023 contains 127 events from 39 repeaters  [^1] In fact, CHIME/FRB Collaboration et al. (2023) marked 14 FRB sources as repeater candidates due to lower significance, indicating that the burst-to-burst DM and sky position differences are larger compared to other confirmed repeaters, and their repetition rates are relatively low. In this study, we consider these 14 FRB sources as repeaters, as they warrant further follow-up observations for potential confirmation., or 151 sub-bursts for all events. Since there are 6 FRBs in Cat1 with flux and fluence values of zero, we excluded those FRBs. Additionally, 6 non-repeaters in Cat1, identified as repeaters, are duplicates in Cat2023, resulting in 739 FRB bursts used in this paper.
@@ 2.2 Feature Selection
Based on these data, we choose the parameters of the sub-bursts as the features for machine learning algorithms. To provide a more comprehensive description of an FRB event, we selected as many parameters as possible to serve as features. These parameters can be divided into two categories: observational parameters (original data provided by Cat1 and Cat2023) and derived parameters, and we present the distribution of all parameters for the entire FRB dataset in Figure 1. We chose 10 observational parameters as did by Chen et al. (2022):
  • Boxcar width $\Delta t_{bc}$ (ms) – The boxcar width of sub-burst, with the label name ‘bc_width’ in two catalogs.
  • Width of sub-burst $\Delta t_{fitb}$ (ms) – The width of sub-burst that is fitted by fitburst [^2] https://github.com/CHIMEFRB/fitburst, with the label name ‘width_fitb’ in two catalogs.
  • Flux $S_{\nu}$ (Jy) – The peak flux of the band-average profile (lower limit) with the label name ‘flux’ in two catalogs.
  • Fluence $F_{\nu}$ ($\rm Jy\cdot ms$) – The flux integrated over the duration of sub-burst (lower limit) with the label name ‘fluence’ in two catalogs.
  • Scattering time $\Delta t_{sc}$ (ms) – The scattering time at 600 MHz of each sub-burst, with the label name ‘scat_time’ in two catalogs.
  • Spectral index $\gamma$ – The spectral shape parameter of sub-burst. The label name in two catalogs is ‘sp_idx’.
@@
  • Highest frequency $\nu_{max}$ (MHz) – The highest frequency band of detection for the sub-burst at full-width tenth-maximum. The label name in two catalogs is ‘high_freq’.
  • Lowest frequency $\nu_{min}$ (MHz) – The lowest frequency band of detection for the sub-burst at full-width tenth-maximum. The label name in two catalogs is ‘low_freq’.
  • Peak frequency $\nu_{p}$ (MHz) – The peak frequency for the sub-burst and is labeled as ‘peak_freq’ in both catalogs.
For $\Delta t_{bc}$, $\Delta t_{fitb}$, $S_{\nu}$, $F_{\nu}$, and $\Delta t_{sc}$, we take their logarithmic values throughout this work. Cat1 and Cat2023 only provide upper limits for the width of sub-burst and scattering time for some FRBs, so we opted to use these upper limits in our analysis.
@@
    The FRB can be separated into different components (see e.g. Deng & Zhang, 2014; Gao et al., 2014; Zhou et al., 2014; Yang et al., 2017; Yang & Zhang, 2016; Li et al., 2019; Wei et al., 2019; Qiang et al., 2020; Qiang & Wei, 2020, 2021; Qiang et al., 2022):
@@
    where $\rm DM_{MW},DM_{halo},DM_{IGM}$ and $\rm DM_{host}$ represent the contributions from the Milky Way, the Milky Way halo, the intergalactic medium (IGM), and the host galaxy (including interstellar medium of the host galaxy and the plasma around source) of the FRB. In the literature, we usually use extragalactic DM for research
@@
    where $c$ is the speed of light, $G$ is the gravitational constant, $m_{p}$ is the mass of proton, $E(z)$ is the dimensionless Hubble parameter. For cosmological parameters, we adopt $H_{0}=67.4$ $\rm km\hskip 2.40005pts^{-1}\hskip 2.40005ptMpc^{-1}$ , $\Omega_{m}=0.315$ and $\Omega_{b,0}h^{2}=0.0224$ from latest Planck 18 results for the flat $\Lambda$CDM cosmology (Planck Collaboration et al., 2020). $f_{e}(z)$ is the ionized electron number fraction per baryon, and $f_{\rm IGM}(z)$ is the fraction of baryon mass in IGM. In principle, these two parameters are functions of redshift z. In this work, we follow e.g. Qiang et al. (2020); Qiang & Wei (2021); Gao et al. (2014); Yang et al. (2017) and use the fiducial values $f_{e}=0.875$ and $f_{\rm IGM}=0.83$. According to Eq.2 and Eq. 3, we can derive the redshift of all FRBs. Following Zhu-Ge et al. (2023), we also set minimum redshift of 0.002248 corresponding to a luminosity distance of 10 Mpc to avoid zero or negative values.
@@
  • Rest-frame width $\Delta t_{rw}$ (ms) – The width of sub-burst $\Delta t_{fitb}$corrected for the cosmological redshift effect: $\Delta t_{rw}$ = $\Delta t_{fitb}$$/(1+z)$. We take the logarithmic values.
  • Burst energy $E$ (erg) – The energy of FRB can be calculated by
@@
  • Brightness temperature $T_{B}$ (K) – The brightness temperature can be derived as
    $T_{B}=\frac{S_{\nu}d_{L}^{2}}{2\pi\kappa_{B}(\nu_{p}\Delta t_{fitb})^{2}}=1.1\times 10^{35}\,{\rm K}\,\left(\frac{S_{\nu}}{\rm Jy}\right)\,\left(\frac{d_{L}}{\rm Gpc}\right)\,\left(\frac{\nu_{p}}{\rm GHz}\right)^{-2}\,\left(\frac{\Delta t_{fitb}}{\rm ms}\right)^{-2},$ (7)

### v2

@@ Abstract
Fast Radio Bursts (FRBs) are highly energetic millisecond-duration astrophysical phenomena typically categorized as repeaters or non-repeaters. However, observational limitations may result in misclassifications, potentially leading to a higher proportion of repeaters than currently identified. In this study, we leverage unsupervised machine learning techniques to classify FRBs using data from the CHIME/FRB catalogs, including both the first catalog and a recent repeater catalog. By employing Uniform Manifold Approximation and Projection (UMAP) for dimensionality reduction and clustering algorithms (k-means and Hierarchical Density-Based Spatial Clustering of Applications with Noise [HDBSCAN]), we successfully segregate repeaters and non-repeaters into distinct clusters, identifying over 100 potential repeater candidates. Our analysis reveals several empirical relations within the clusters, including the ${\rm log\,}\Delta t_{sc}-{\rm log\,}\Delta t_{rw}$, ${\rm log\,}\Delta t_{sc}-{\rm log\,}T_{B}$, and $r-\gamma$ correlations, where ${\Delta t_{sc},\Delta t_{rw},T_{B},r,\gamma}$ represent scattering time, rest-frame width, brightness temperature, spectral running, and spectral index, respectively. The Chow test results reveal that while some repeaters and non-repeaters share similar empirical relationships, the overall distinctions between the two groups remain significant, reinforcing the classification of FRBs into repeaters and non-repeaters. These findings provide new insights into the physical properties and emission mechanisms of FRBs. This study demonstrates the effectiveness of unsupervised learning in classifying FRBs and identifying potential repeaters, paving the way for more precise investigations into their origins and applications in cosmology. Future improvements in observational data and machine learning methodologies are expected to further enhance our understanding of FRBs.
@@ 1 Introduction
To explore their origin, numerous radio telescopes have dedicated time and efforts on FRBs, such as, the Deep Synoptic Array (DSA; Kocz et al., 2019; Hallinan et al., 2019), Arecibo (Spitler et al., 2014), Parkes (e.g., Lorimer et al., 2007; Burke-Spolaor & Bannister, 2014; Petroff et al., 2015; Ravi & Lasky, 2014), the Canadian Hydrogen Intensity Mapping Experiment (CHIME; CHIME/FRB Collaboration et al., 2018), the Five-hundred-meter Aperture Spherical Radio Telescope (FAST; Li & Pan, 2016), and the Australian Square Kilometer Array Pathfinder (ASKAP; Shannon et al., 2024). So far, nearly a thousand FRBs have been observed (Petroff et al., 2016; CHIME/FRB Collaboration et al., 2021; Jankowski et al., 2023; Xu et al., 2023), with nearly 100 of them have known redshifts (e.g. Law et al., 2024; Sharma et al., 2024; Bhardwaj et al., 2024; Gordon et al., 2023). To date, the largest FRB sample is the first CHIME/FRB catalog (CHIME/FRB Collaboration et al., 2021).
@@
Based on the observational characteristics, FRBs are broadly classified into two categories, i.e., repeaters and non-repeaters. As the names suggest, repeaters are sources that have exhibited multiple bursts, whereas non-repeaters have only been observed to burst once. Since the discovery of the first repeater, FRB 20121102 (Spitler et al., 2016), more than 50 FRBs have been identified as repeaters to date (CHIME/FRB Collaboration et al., 2021, 2023; Kumar et al., 2019; Kirsten et al., 2022; Fonseca et al., 2020; Niu et al., 2022; Xu et al., 2022), while most of the remaining ones are classified as non-repeaters. However, Ravi (2019) suggests that the volumetric rate of non-repeating FRBs might exceed that of cataclysmic events and the formation rate of compact objects, implying that the majority of FRBs should be repeaters. Some studies also propose that more than half of the FRBs in the first CHIME/FRB catalog could be repeaters (Yamasaki et al., 2023; McGregor & Lorimer, 2024). Furthermore, several FRBs initially identified as non-repeaters were later observed to repeat (CHIME/FRB Collaboration et al., 2023). These suggest that many of the FRBs currently classified as non-repeaters might actually be potential repeaters, with only a single burst detected due to various observational factors. As a result, some studies have attempted to identify potential repeaters among apparent non-repeating FRBs, and one of the methods being employed is machine learning.
@@
Machine learning is an artificial intelligence technique that allows computers to learn from data and make predictions or decisions without being explicitly programmed (Cover & Hart, 1967; Rumelhart et al., 1986). Machine learning is generally categorized into three types, i.e., supervised learning, unsupervised learning, and semi-supervised learning (e.g. Dempster et al., 2018; Breiman, 2001; Chang & Lin, 2011; Vapnik, 1999). Using algorithms to uncover patterns in the data can be applied to tasks such as classification, regression, clustering, and optimization. Currently, machine learning has already been widely used in the detection and analysis of FRBs (e.g. Wagstaff et al., 2016; Zhang et al., 2018; Wu et al., 2019; Yang et al., 2021; Adámek & Armour, 2020; Agarwal et al., 2020; Bhatporia et al., 2023). For instance within the first CHIME/FRB catalog dataset, Chen et al. (2021) and Zhu-Ge et al. (2022) identify 188 and 117 repeater candidates from 474 non-repeating FRBs via unsupervised learning, correspondingly. Yang et al. (2023) instead discovered 145 repeater using FRB morphology as features, while Luo et al. (2022) identified dozens of repeater candidates with various supervised learning methods. Furthermore, Luo et al. (2022) found that the most prominent factors to distinguish between non-repeating and repeating FRBs are brightness temperature and rest-frame frequency bandwidth, whereas Sun et al. (2024) found spectral running may play a role instead. Furthermore, some studies have used machine learning techniques to classify thousands of bursts from highly active repeaters, such as FRB 20121102 (Raquel et al., 2023) and FRB 20201124A (Chen et al., 2023), in an effort to analyze their potential radiation mechanisms.
@@
In addition to classifying FRBs based on their repeatability, some studies have also investigated the clustering of FRBs using characteristics beyond repeatability, as well as the empirical relationships among these features. For instance, similar to the well-known classification of GRB, where short GRBs exhibit short prompt emissions and are mostly associated with the old population, while long GRBs with long prompt emissions and are mostly associated with the young population (Zhang et al., 2007; Kumar & Zhang, 2014), Guo & Wei (2022) proposed classifying FRBs based on their association with either old or young stellar populations. They discovered several tight empirical relations for non-repeaters in the first CHIME/FRB catalog, such as ${\rm log\,}E-{\rm log\,}L_{\nu}$, ${\rm log\,}E-{\rm log\,DM_{E}}$ and ${\rm log\,DM_{E}}-{\rm log\,}L_{\nu}$, where ${\rm log\,}$ means the logarithm to base 10, and $E$, ${\rm DM_{E}}$, $L_{\nu}$ are isotropic energy, spectral luminosity, and extragalactic DM, respectively. Similar empirical relations were found for non-repeaters associated with old populations and all non-repeaters, though with notably different slopes and intercepts, such as ${\rm log\,}F_{\nu}-{\rm log\,}S_{\nu}$, ${\rm log\,DM_{E}}-{\rm log\,}S_{\nu}$, and ${\rm log\,}F_{\nu}-{\rm log\,DM_{E}}$, where $F_{\nu}$ is specific fluence, $S_{\nu}$ is flux. Many empirical relations still hold for localized FRBs (Li et al., 2024). Based on these empirical relations, FRBs could potentially serve as standard candles, allowing cosmological models to be constrained without relying on DM measurements (Guo & Wei, 2024). Li et al. (2021) classified FRBs into short ($W<100\,$ms) and long ($W>100\,$ms) bursts, where $W$ represents the pulse width. A strong power-law correlation between fluence and peak flux density was identified for these categories. Xiao & Dai (2022) classified repeating FRBs into classical ($T_{B}\geq 10^{33}\,$) and atypical ($T_{B}<10^{33}\,$) bursts, where $T_{B}$ refers to the brightness temperature. A tight power-law correlation between pulse width and fluence was also observed for classical bursts.
@@
CHIME has recently released a new repeater catalog (CHIME/FRB Collaboration et al., 2023), significantly increasing the data available on repeaters. This expanded dataset is expected to improve the accuracy of identifying repeater candidates through machine learning. In this paper, we will apply unsupervised learning methods to classify the FRBs from the combined data of these two catalogs into different clusters, find potential repeaters among the non-repeaters, and analyze possible empirical relations across these clusters. This paper is structured as follows. In Section 2, we present the selected CHIME data and the features used in our analysis. Section 3 introduces two types of unsupervised machine learning methods, i.e., dimensionality reduction and clustering, as well as evaluation metrics. In Section 4, we present the results of dimensionality reduction and clustering, then we analyze the potential empirical relations. Finally, Section 5 provides our conclusions.
@@ 2.1 Sample Construction
In this paper, we use the largest available FRB data from CHIME to date, including the first CHIME/FRB Catalog (CHIME/FRB Collaboration et al., 2021, hereafter Cat1) and the CHIME/FRB Collaboration (2023) Catalog (CHIME/FRB Collaboration et al., 2023, hereafter Cat2023). The Cat1 includes 536 events (474 non-repeaters and 62 repeat bursts from 18 repeaters). There are 600 sub-bursts in total because of multiple peaks appearing in the light curve of FRB. The Cat2023 contains 127 events from 39 repeaters  [^1] In fact, CHIME/FRB Collaboration et al. (2023) marked 14 FRB sources as repeater candidates due to lower significance, indicating that the burst-to-burst DM and sky position differences are larger compared to other confirmed repeaters, and their repetition rates are relatively low. In this study, we consider these 14 FRB sources as repeaters, as they warrant further follow-up observations for potential confirmation., or 151 sub-bursts for all events. Since there are 6 FRBs in Cat1 with flux and fluence values of zero, we excluded those FRBs. Additionally, 6 non-repeaters in Cat1, identified as repeaters, are duplicates in Cat2023, resulting in 739 FRB bursts used in this paper.
@@ 2.2 Feature Selection
We treat all sub-bursts as individual bursts and use their features for machine learning algorithms. To provide a more comprehensive description of an FRB event, we included a broad range of parameters based on their usage in previous research, as different studies emphasize different features and identify varying important ones (e.g. Chen et al., 2022; Luo et al., 2022; Zhu-Ge et al., 2022; Sun et al., 2024). These parameters can be divided into two categories: observational parameters (original data provided by Cat1 and Cat2023) and derived parameters, and we present the distribution of all parameters for the entire FRB dataset in Figure 1. We chose 10 observational parameters as did by Chen et al. (2022):
  • Boxcar width $\Delta t_{bc}$ (ms) – The boxcar width of the burst, with the label name ‘bc_width’ in the two catalogs.
  • Width of the burst $\Delta t_{fitb}$ (ms) – The width of the burst that is fitted by fitburst [^2] https://github.com/CHIMEFRB/fitburst, with the label name ‘width_fitb’ in the two catalogs.
  • Flux $S_{\nu}$ (Jy) – The peak flux of the band-average profile (lower limit) with the label name ‘flux’ in the two catalogs.
  • Fluence $F_{\nu}$ ($\rm Jy\cdot ms$) – The flux integrated over the duration of the burst (lower limit) with the label name ‘fluence’ in the two catalogs.
  • Scattering time $\Delta t_{sc}$ (ms) – The scattering time at 600 MHz of the burst, with the label name ‘scat_time’ in the two catalogs.
  • Spectral index $\gamma$ – The spectral shape parameter of the burst. The label name in the two catalogs is ‘sp_idx’.
@@
  • Highest frequency $\nu_{max}$ (MHz) – The highest frequency band of detection for the burst at full-width tenth-maximum. The label name in the two catalogs is ‘high_freq’.
  • Lowest frequency $\nu_{min}$ (MHz) – The lowest frequency band of detection for the burst at full-width tenth-maximum. The label name in the two catalogs is ‘low_freq’.
  • Peak frequency $\nu_{p}$ (MHz) – The peak frequency for the burst and is labeled as ‘peak_freq’ in both catalogs.
For $\Delta t_{bc}$, $\Delta t_{fitb}$, $S_{\nu}$, $F_{\nu}$, and $\Delta t_{sc}$, we take their logarithmic values throughout this work. Cat1 and Cat2023 only provide upper limits for the width of the burst and scattering time for some FRBs, so we opted to use these upper limits in our analysis.
@@
    In astronomical observations, distance plays a crucial role in analyzing the origin of FRBs. The Dispersion Measure (DM), indicating the total column density of free electrons along the line of sight, serves as a key distance-related parameter. For most FRBs, DM values far exceed those predicted for the Milky Way, pointing to their extragalactic origins. The DM of FRB can be separated into different components (see e.g. Deng & Zhang, 2014; Gao et al., 2014; Zhou et al., 2014; Yang et al., 2017; Yang & Zhang, 2016; Li et al., 2019; Wei et al., 2019; Qiang et al., 2020; Qiang & Wei, 2020, 2021; Qiang et al., 2022):
@@
    where $\rm DM_{MW},DM_{halo},DM_{IGM}$ and $\rm DM_{host}$ represent the contributions from the Milky Way, the Milky Way halo, the intergalactic medium (IGM), and the host galaxy (including the interstellar medium of the host galaxy and the plasma around source) of the FRB. In the literature, we usually use extragalactic DM for research
@@
    where $c$ is the speed of light, $G$ is the gravitational constant, $m_{p}$ is the mass of proton, $E(z)$ is the dimensionless Hubble parameter. For cosmological parameters, we adopt $H_{0}=67.4$ $\rm km\hskip 2.40005pts^{-1}\hskip 2.40005ptMpc^{-1}$ , $\Omega_{m}=0.315$ and $\Omega_{b,0}h^{2}=0.0224$ from the latest Planck 18 results for the flat $\Lambda$CDM cosmology (Planck Collaboration et al., 2020). $f_{e}(z)$ is the ionized electron number fraction per baryon, and $f_{\rm IGM}(z)$ is the fraction of baryon mass in IGM. In principle, these two parameters are functions of redshift z. In this work, we follow e.g. Qiang et al. (2020); Qiang & Wei (2021); Gao et al. (2014); Yang et al. (2017) and use the fiducial values $f_{e}=0.875$ and $f_{\rm IGM}=0.83$. According to Eq.2 and Eq. 3, we can derive the redshift of all FRBs. Following Zhu-Ge et al. (2023), we also set a minimum redshift of 0.002248 corresponding to a luminosity distance of 10 Mpc to avoid zero or negative values.
@@
  • Rest-frame width $\Delta t_{rw}$ (ms) – The width of the burst $\Delta t_{fitb}$ corrected for the cosmological redshift effect: $\Delta t_{rw}$ = $\Delta t_{fitb}$$/(1+z)$. We take the logarithmic values.
  • Burst energy $E$ (erg) – The energy of a FRB can be calculated by
@@
  • Brightness temperature $T_{B}$ (K) – The brightness temperature can be derived as (Luo et al., 2022)
    $T_{B}=\frac{S_{\nu}d_{L}^{2}}{2\pi\kappa_{B}(\nu_{p}\Delta t_{fitb})^{2}}=1.1\times 10^{35}\,{\rm K}\,\left(\frac{S_{\nu}}{\rm Jy}\right)\,\left(\frac{d_{L}}{\rm Gpc}\right)^{2}\,\left(\frac{\nu_{p}}{\rm GHz}\right)^{-2}\,\left(\frac{\Delta t_{fitb}}{\rm ms}\right)^{-2}\,\frac{1}{1+z},$ (7)